At $23 per square yard, what would carpet cost for a room that measures 36 feet by 12 feet?
: At $23 per square yard, what would carpet cost for a room that measures 36 feet by 12 feet?

Answer by checkley77(3848)

36*12=432 ft^2
[432/9=48] yd^2 [forgot to convert ft^2 to yd^2]
48*23=$1,104.00 cost of the carpet.